OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 2y = 12?

OpenStudy (saifoo.khan):

-3/2 slope

OpenStudy (saifoo.khan):

y intercpet when x = 0 x =4

OpenStudy (asadkarim7):

y = -3/2 x +6 m = -3/2 y intercept 6

OpenStudy (saifoo.khan):

asad, im going! Allah hafiZz!